# Big Reasoning with Small Models: Instruction Retrieval at Inference Time

### v1

@@ 4 Experimental Setup
We evaluate the effect of the Instruction corpus on three benchmarks and publicly available models spanning five major open families, and compare performance against GPT-4o as a reference frontier model.
@@ 4.1 Tasks
We use three benchmarks that collectively stress different dimensions of reasoning. MedQA (Jin et al., 2020) contains multiple-choice questions from professional medical board exams; success requires both factual recall of biomedical knowledge and diagnostic reasoning. MMLU Professional Law (Hendrycks et al., 2021a) focuses on legal exam questions, testing recall of statutes and precedents alongside structured case-based reasoning. MathQA (Amini et al., 2019) evaluates symbolic reasoning through math word problems, requiring models to translate natural language into computational procedures.
@@ 4.2 Models
We evaluate across thirteen models from five major open families: Llama 3 (AI@Meta, 2024), Gemma 2 (Team, 2024), Qwen 3 (Team, 2025), Mistral (Jiang et al., 2023), and DeepSeek R1 Distilled (DeepSeek-AI, 2025). Model sizes span from 1B to 14B parameters, providing coverage across the typical range of deployable SLMs. For Llama 3, Gemma 2, and Mistral we use publicly released instruction-tuned variants. For DeepSeek R1 we evaluate distilled models released in sizes from 1.5B to 14B. For Qwen 3, we include both base models (1.7B, 4B, 14B) and an instruction-tuned variant (4B-Instruct). This mixture allows us to compare instruction-following capabilities directly against base model behavior. No additional fine-tuning is applied. At inference, each test question is embedded using the same model employed during clustering (text-embedding-3-large) and matched to training clusters by cosine similarity. The top five cluster-level instructions are retrieved and prepended to the query before decoding. To ensure comparability, all prompts and inference settings were standardized across tasks, with generation performed at a temperature of 0.7 and a top-p value of 0.95. GPT-4o, accessed via API, is included as a reference point to situate results against current frontier performance.
@@ 5.1 Aggregate Performance
Figure 2 shows the accuracy of the High School Concise instructions compared to zero-shot prompting across MathQA, MedQA, and MMLU Law. For the smallest models under 3B parameters, improvements are mixed and often negative. A few exceptions exist, such as Qwen-3 1.7B base, which gains +15pp on MedQA, but the same model shows negligible changes on MMLU and MathQA. This instability suggests that tiny models lack the intrinsic reasoning ability to reliably leverage retrieved scaffolds, which is in line with prior research Li et al. (2025). At around 3B parameters, gains become consistently positive, even if modest. LLaMA-3 3B, for instance, improves by +1–5pp across benchmarks. This marks a capacity threshold: once a model is large enough to follow multi-step guidance, retrieved instructions shift from being noise to providing usable structure. Above this threshold, performance consistently improves. Models from 4B to 14B shows gains across all three tasks, with deltas typically in the +5–18pp range. The largest improvements appear in MedQA and MMLU Law, where domain-specific knowledge and reasoning may lack representation in the limited parameter space. For example, Gemma-2 9B improves by +14pp on MMLU Law and Mistral 7B by +16pp, while MedQA gains reach +18pp at 9B. In contrast, MathQA gains are more muted, though still positive, likely because mathematical reasoning is already well represented in pretraining data, leaving less headroom for improvement. Overall, retrieved instructions consistently boost accuracy once models exceed 3B parameters, with robust gains across both symbolic and domain-expert reasoning tasks. Notably, on MedQA and MMLU Law, the 14B parameter SLMs with retrieved instructions surpass the zero-shot accuracy of GPT-4o.
@@ 6 Ablation and Analysis
While instruction retrieval improves small-model reasoning, the source of these gains remains unclear. We perform targeted ablations on MedQA using the High School Concise variant and models of at least 3B parameters to identify what drives improvement.
6.1 Component Ablations
Do instruction gains stem from retrieved background knowledge, reasoning steps, or their combination? To isolate these effects, we ablate the two components of each instruction on MedQA. In the knowledge-only condition, instructions include only the background section, providing factual context but no reasoning steps. In the reasoning-only condition, instructions include only the procedural steps, with no supporting knowledge. The full condition retains both.
Figure 5: Performance impact of ablating instruction components on MedQA (High School Concise). Both knowledge-only and reasoning-only variants outperform zero-shot prompting, though the relative advantage differs by model family.
Both components independently improve performance over the zero-shot baseline (Figure 5). On average, removing either section reduces accuracy relative to the full instruction, confirming that background knowledge and procedural steps contribute complementary information. Across models, the knowledge-only variant outperforms the reasoning-only variant by roughly two percentage points on average, suggesting that factual grounding plays a slightly stronger role in MedQA. The pattern, however, varies by model family: Qwen and DeepSeek models benefit most from factual context, whereas Gemma gains more from reasoning steps. In short, impact varies with each model’s strengths: models lacking domain knowledge rely more on factual context, whereas those with limited reasoning ability benefit more from explicit steps.
@@ 7 Discussion
Our findings show that small models above a minimal capacity threshold can reliably incorporate retrieved instructions to perform multi-step reasoning, narrowing the gap to large models without additional training. This pattern aligns with prior evidence (Li et al., 2025) that models under roughly 3B parameters struggle to internalize long reasoning traces but can effectively leverage shorter, externally provided guidance. By treating instructions as modular, retrievable assets, our approach shifts part of the reasoning process from model parameters to structured text, creating a practical foundation for more maintainable and auditable small models. Externalized instructions are non-parametric and can be updated or versioned without retraining, enabling faster and safer revisions in regulated domains. Because they are written in natural language, instructions are also interpretable and portable across model families, allowing a single high-quality corpus to provide consistent reasoning support for diverse architectures. In applied settings, this separation makes reasoning assistance feasible in privacy-constrained environments: for example, a 7B-parameter model running locally on a clinician’s laptop can retrieve targeted instructions for diagnostic reasoning or documentation without transmitting patient data. Ultimately, instruction effectiveness depends on a model’s existing competence and confidence. Domains where models already exhibit strong internal reasoning, such as mathematics, show smaller gains from retrieval because the external instructions provide limited additional value. The same principle likely applies at the instance level: even within a single model, some queries require external guidance while others can be solved unaided. Future systems should therefore adapt retrieval based on a model’s estimated capability and uncertainty, invoking instructions only when they are expected to improve reasoning. Developing such competence-aware or confidence-gated retrieval mechanisms is a key step toward making instruction intervention dependable.
Figure 6: Effect of clustering granularity on MedQA accuracy, averaged across all evaluated models. Performance remains consistent across thresholds, indicating that instruction retrieval is robust even when clusters merge into broader groups.
@@ 8 Conclusion
This work reframes reasoning as a retrieval problem: rather than generating reasoning chains from scratch, small language models retrieve structured instructions that integrate background knowledge with step-by-step reasoning to guide their generation. Across MedQA, MMLU Law, and MathQA, this approach consistently improves multi-step reasoning without additional training. Concise, well-structured instructions yield the largest gains, and improvement depends more on a model’s ability to follow structure than on size alone. These findings suggest that both factual knowledge and reasoning procedures can be externalized as reusable text modules, enabling small models to approximate large-model performance while preserving efficiency and privacy. Looking ahead, competence-aware retrieval mechanisms that adapt instruction use based on model confidence may further bridge the gap between compact and frontier-scale reasoning systems.
@@ A.1.4 Variant Evaluation Results
Graduate Concise 4.57 4.92 4.83 4.90 4.95
Graduate Long 4.97 4.69 4.93 4.98 4.94
High school Concise 4.31 4.91 4.73 4.81 4.91
High school Long 4.68 4.82 4.90 4.95 4.91

### v2

@@ 4 Experimental Setup
We evaluate the effect of the Instruction corpus on three benchmarks and publicly available models spanning five major open families, and compare performance against GPT-4o as a reference LLM.
@@ 4.1 Tasks
We use three benchmarks that collectively stress different dimensions of reasoning. MedQA (Jin et al., 2020) contains multiple-choice questions from professional medical board exams; success requires both factual recall of biomedical knowledge and diagnostic reasoning. MMLU Professional Law (Hendrycks et al., 2021a) focuses on legal exam questions, testing recall of statutes and precedents alongside structured case-based reasoning. MathQA (Amini et al., 2019) evaluates symbolic reasoning through math word problems, requiring models to translate natural language into computational procedures. For MedQA, we additionally compare against a standard RAG baseline using the benchmark’s provided medical textbooks. The other two tasks do not allow for a comparable RAG baseline without constructing a new external corpus. Given a test question, the model retrieves the top-$5$ relevant passages using the same embedding model (text-embedding-3-large) and cosine similarity; the retrieved passages are provided in the prompt at inference time.
@@ 4.2 Models
We evaluate across thirteen models from five major open families: Llama 3 (AI@Meta, 2024), Gemma 2 (Team, 2024), Qwen 3 (Team, 2025), Mistral (Jiang et al., 2023), and DeepSeek R1 Distilled (DeepSeek-AI, 2025). Model sizes span from 1B to 14B parameters, providing coverage across the typical range of deployable SLMs. For Llama 3, Gemma 2, and Mistral we use publicly released instruction-tuned variants. For DeepSeek R1 we evaluate distilled models released in sizes from 1.5B to 14B. For Qwen 3, we evaluate three model sizes (1.7B, 4B, and 14B), using the default thinking-enabled configuration. At inference, each test question is embedded using the same model used during clustering (text-embedding-3-large) and matched to training clusters by cosine similarity. The top five cluster-level instructions are retrieved and inserted into the prompt before decoding. To ensure comparability, all prompts and inference settings were standardized across tasks, with generation performed at a temperature of 0.7 and a top-p value of 0.95. GPT-4o, accessed via API, is included as a reference point to situate results against LLM performance.
@@ 5.1 Aggregate Performance
Figure 2 shows the accuracy of the High School Concise instructions compared to zero-shot prompting across MathQA, MedQA, and MMLU Law. For the smallest models under 3B parameters, improvements are mixed and often negative. A few exceptions exist, such as Qwen-3 1.7B base, which gains +15pp on MedQA, but the same model shows negligible changes on MMLU and MathQA. This instability suggests that tiny models lack the intrinsic reasoning ability to reliably leverage retrieved scaffolds, which is in line with prior research Li et al. (2025). At around 3B parameters, gains become consistently positive, even if modest. LLaMA-3 3B, for instance, improves by +1–5pp across benchmarks. Once a model is large enough to follow multi-step guidance, retrieved instructions shift from being noise to providing usable structure and information. Notably, both instruction retrieval and passage-based RAG reduce accuracy for models below 3B parameters. Once models cross this threshold, instruction retrieval yields substantially larger gains than RAG on MedQA (+9.3pp vs. +3.2pp on average), suggesting that structured procedural guidance is more effective than unstructured evidence for small-model reasoning. Above this threshold, performance consistently improves for instructions. Models from 4B to 14B shows gains across all three tasks, with deltas typically in the +5–18pp range. The largest improvements appear in MedQA and MMLU Law, where domain-specific knowledge and reasoning may lack representation in the limited parameter space. For example, Gemma-2 9B improves by +14pp on MMLU Law and Mistral 7B by +16pp, while MedQA gains reach +18pp at 9B. In contrast, MathQA gains are more muted, though still positive, likely because mathematical reasoning is already well represented in pretraining data, leaving less headroom for improvement. Overall, retrieved instructions consistently boost accuracy once models exceed 3B parameters, with robust gains across both symbolic and domain-expert reasoning tasks. Importantly, on MedQA and MMLU Law, the 14B parameter SLMs with retrieved instructions surpass the zero-shot accuracy of GPT-4o.
@@ 6 Ablation and Analysis
While instruction retrieval consistently improves small-model reasoning, its effectiveness depends on design choices such as corpus granularity and construction cost. We analyze these factors on MedQA using the High School Concise variant and models of at least 3B parameters, focusing on robustness to clustering choices and the amortized cost of corpus construction.
@@ 6.1 Cluster Size
6.2 Amortized Cost Analysis
Instruction retrieval introduces a one-time processing cost to construct the Instruction Corpus, after which inference proceeds with a fixed retrieval overhead. This design shifts computation from repeated generation to amortized preparation, making the approach cost-effective even at moderate corpus sizes. We estimate corpus construction costs using GPT-5.1, assuming an average of 300 input tokens and 600 output tokens per instruction, reflecting our empirical finding that concise instructions perform best. At current pricing (\$1.25 per million input tokens and \$10.00 per million output tokens), generating a single instruction costs approximately \$0.0064. Under these assumptions, a corpus of 1,000 instructions costs \$6.38 to generate, while a corpus of 10,000 instructions costs \$63.75. Embedding costs are negligible in comparison. Using text-embedding-3-large at \$0.065 per million tokens (batched), embedding the entire corpus adds less than \$1, even at the largest corpus sizes. Importantly, our ablation results show that coarser clustering, and thus fewer, more general instructions, perform comparably to highly bespoke corpora. In practice, corpora in the 1k–3k range achieve most of the observed gains, placing total one-time costs well below \$20. Importantly, this cost is incurred only once and can be amortized across an unlimited number of users and inference runs. For example, a hospital system could construct an instruction corpus derived from clinical guidelines for diagnosing heart failure, which could then be reused indefinitely by cardiologists across multiple hospitals without additional generation cost.
@@ 7 Discussion
Our results show that small language models above a minimal capacity threshold can reliably incorporate retrieved instructions to perform multi-step reasoning, substantially narrowing the gap to large models without additional training. This finding aligns with prior evidence (Li et al., 2025) that models below roughly 3B parameters struggle to internalize long reasoning traces, but can benefit from shorter, externally provided guidance. Once models cross this threshold, retrieved instructions shift from being noise to providing usable structure, suggesting that instruction retrieval exploits reasoning capabilities that are present but underutilized in small models. Although instruction retrieval yields consistent gains, the mechanisms underlying these improvements remain an open question. One plausible explanation is that retrieved instructions supply highly targeted domain knowledge that small models cannot reliably recall from their parameters alone. Another is that the explicit procedural structure reduces search during generation by constraining the order and priority of reasoning steps, leading to more stable and coherent inference. In practice, these effects are likely complementary: instructions both surface relevant information and impose a reasoning policy that small models can execute more reliably than unconstrained chain-of-thought generation.
Our results also suggest that instruction retrieval should not be applied indiscriminately. Gains are largest on knowledge-intensive domains such as medicine and law, where both factual recall and structured decision-making exceed the capacity of small models. In contrast, improvements on mathematics are smaller, reflecting limited headroom in domains where models already exhibit strong internal reasoning. These findings motivate several directions for future work. In the current framework, instructions are retrieved for every query, even when a model may already be capable of solving the problem unaided. A natural extension is competence-aware retrieval, in which models estimate their confidence or uncertainty and invoke instructions only when external guidance is likely to improve reasoning. More broadly, because instructions are external, non-parametric artifacts, instruction corpora need not be static. Future systems could support self-evolving instruction collections that update over time as models encounter new failure modes, domain knowledge changes, or revised best practices, enabling continuous improvement without retraining.
Figure 5: Effect of clustering granularity on MedQA accuracy, averaged across all evaluated models. Performance remains consistent across thresholds, indicating that instruction retrieval is robust even when clusters merge into broader groups. Broken down by model size in Appendix 8
@@ 8 Conclusion
This work reframes reasoning as a retrieval problem. Rather than requiring small language models to internally generate or store specialized domain knowledge or complex reasoning chains, we show that they can retrieve and execute structured instructions that combine domain knowledge with explicit procedures. Across MedQA, MMLU Law, and MathQA, instruction retrieval consistently improves multi-step reasoning for models above a minimal capacity threshold, without any additional training. Concise instructions yield the largest gains, and effectiveness depends more on a model’s ability to follow structured guidance than on parameter count alone. Unlike chain-of-thought prompting, which relies on scale, or distillation, which requires task-specific training and often reduces generality, instruction retrieval externalizes reasoning as reusable text. Compared to standard retrieval-augmented generation, which provides unstructured information, instructions define an explicit and inspectable problem-solving strategy. By decoupling reasoning knowledge from model parameters, instruction retrieval enables small models to approach large-model performance while preserving the efficiency and privacy of local inference. A single instruction corpus can be reused and updated independently of the model, offering a practical and maintainable path toward reliable reasoning with compact models.
@@ A.1.4 Variant Evaluation Results
Graduate Concise 4.95 4.84 4.99 5.00 4.87
Graduate Long 5.00 4.53 5.00 4.99 4.48
High school Concise 4.64 4.83 4.97 4.99 4.89
High school Long 4.99 4.76 4.99 5.00 4.88
@@ Appendix C Ablation
Figure 8: MedQA ablation results by model size. Instruction retrieval is relatively robust to clustering granularity for models $\geq$3B parameters. Brokenn down by model size
@@
Model Zero-Shot RAG Instructions RAG $\Delta$ Instr $\Delta$
LLaMA-3 1B Instruct 0.30 0.22 0.19 -0.08 -0.11
DeepSeek-R1 Qwen 1.5B 0.21 0.26 0.31 +0.05 +0.10
Qwen-3 1.7B Base 0.45 0.53 0.60 +0.07 +0.15
Gemma-2 2B IT 0.36 0.27 0.32 -0.09 -0.04
LLaMA-3 3B Instruct 0.49 0.48 0.50 -0.01 +0.01
Qwen-3 4B Base 0.64 0.68 0.73 +0.04 +0.08
Qwen-3 4B Instruct 0.64 0.69 0.74 +0.05 +0.10
DeepSeek-R1 Qwen 7B 0.34 0.37 0.47 +0.03 +0.12
Mistral 7B Instruct 0.33 0.38 0.44 +0.05 +0.11
DeepSeek-R1 LLaMA 8B 0.51 0.56 0.60 +0.05 +0.09
Gemma-2 9B IT 0.37 0.40 0.54 +0.03 +0.18
DeepSeek-R1 Qwen 14B 0.68 0.69 0.74 +0.01 +0.06
Qwen-3 14B Base 0.73 0.78 0.83 +0.04 +0.09
Table 7: MedQA: Zero-Shot vs RAG (Top-5 Passages) vs Instructions. RAG retrieval from medical textbooks provides modest gains (+2 $p$ avg), while instruction retrieval achieves larger improvements (+7 $p$ avg). Both methods hurt performance on small models (1-2B parameters). $\Delta$ denotes the difference from zero-shot accuracy.
@@ Appendix D Instruction Prompt Templates
# Question
{QUESTION}
# Required Output Format
‘‘‘json
"reasoning": "<step-by-step analysis of the medical scenario and answer choices>",
"final_answer": "A"
Figure 9: Zero-shot Prompt
@@
Question
{QUESTION}
# Retrieved Context
The following passages from medical textbooks may be relevant to answering this question.
{PASSAGES}
# Required Output Format
‘‘‘json
"reasoning": "<step-by-step analysis of the medical scenario and answer choices>",
"final_answer": "A"
Figure 15: MedQA RAG baseline Prompt